You want to have $11,000 saved ten years from now. How much less do you have to deposit today to reach this goal if you can earn 7% rather than 6% on your savings?
Answer #1
Amount to be deposited today at 6% = 11000/1.06^10 = $   6,142.34
Amount to be deposited today at 7% = 11000/1.07^10 = $   5,591.84
Reduction in amount to be deposited today at 7%, when compared to the interest rate of 6% = $      550.50